OpenStudy (anonymous):

I will fan and medal! If θ is in Quadrant IV and sinθ=(-5/13), then secθ=?

OpenStudy (midhun.madhu1987):



OpenStudy (anonymous):

cos theta= 12/13 sec theta =13/12

OpenStudy (midhun.madhu1987):

:) :)

OpenStudy (anonymous):

so its going to be positive?

OpenStudy (midhun.madhu1987):

yes...

OpenStudy (anonymous):

okay thank you again @midhun.madhu1987

OpenStudy (midhun.madhu1987):

np